A container has 1 red marble 2 green marbles 3 yellow marbles and 4 orange marbles. You select 2 marbles without replacement.
Find:
---Probability of selecting 2 orange marbles.
--Probability of selecting 2 red marbles.
--Probability of selecting atleast 1 red marble.
---Probability that the second marble is yellow given the first marble is green.

Answer by stanbon(75887)

A container has 1 red marble 2 green marbles 3 yellow marbles and 4 orange marbles. You select 2 marbles without replacement.
Find:
---Probability of selecting 2 orange marbles = 4C2/10C2 = 6/45 = 2/15
------------------------------------------------------
--Probability of selecting 2 red marbles = 0/45 = 0
------------------------------------------------------
--Probability of selecting atleast 1 red marble:::
1 - 9C2/10C2 = 1-36/45 = 9/45 = 1/5
-------------------------------------------------------
---Probability that the second marble is yellow given the first marble is green.
= P(yellow AND green)/P(green) = [(3*2)/10C2]/[2/10] = (6/45)/(1/5) = 2/3
---------------------------------------------------------
Cheers,
Stan H.
=======================